What is the %(m/m) of a solution containing 25.7 g of lithium fluoride (LiF) dissolved in 79.3 g of water?
Answer #1

To calculate %(m/m) (mass percentage) of a solute, the formula used is:

%(m/m) = Mass of solute Total mass of solution - X 100%

Here, solute = LiF , solvent = water

So, mass of solute = 25.7 g

Mass of solution = mass of solute + mass of solvent = 25.7 g + 79.3g = 105 g

So, putting the values, we get:

%(m/m) of LiF = (25.7 / 105) x 100 % = 24.48 %